# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v6

@@ Abstract.
Generative AI has begun to alter how we work, learn, communicate, and participate in online communities. How might our online communities be changed by generative AI? To start addressing this question, we focused on online community moderators’ experiences with AI-generated content (AIGC). We performed fifteen in-depth, semi-structured interviews with moderators of Reddit communities that restrict the use of AIGC. Our study finds that rules about AIGC are motivated by concerns about content quality, social dynamics, and governance challenges. Moderators fear that, without such rules, AIGC threatens to reduce their communities’ utility and social value. We find that, despite the absence of foolproof tools for detecting AIGC, moderators were able to somewhat limit the disruption caused by this new phenomenon by working with their communities to clarify norms. However, moderators found enforcing AIGC restrictions challenging, and had to rely on time-intensive and inaccurate detection heuristics in their efforts. Our results highlight the importance of supporting community autonomy and self-determination in the face of this sudden technological change, and suggest potential design solutions that may help.
@@ 1. Introduction
Recent advances in generative AI (OpenAI22; openai2023gpt4; ramesh2021zeroshot) have led to the proliferation of AI-generated content (AIGC) across the internet (Thompson23Discovered), causing online communities to adjust their policies and practices. Generative AI use, whether disclosed or not, may alter the social dynamics of online communities. For example, AI-generated text with subtle inaccuracies poses a challenge for knowledge-sharing communities, like Stack Exchange (makyen23), and communities where members engage in cooperative work, such as Wikipedia (gertner_wikipedias_2023). In addition, the perceived inauthenticity of AIGC can be disruptive to social communities where people come together for human connection, such as some Facebook groups (fiesler_ai_2024). This disruption adds to the existing challenges faced by online communities, such as coordinated disinformation campaigns, which generative AI may make cheaper, more effective, and better targeted (goldstein23).
@@
As many communities rely on content moderators for stewardship (grimmelmann2015virtues), it is important to understand and document moderators’ firsthand experiences responding to AIGC. Effective moderation can help communities stay civil during times of disruptive change by clarifying and enforcing community norms (Kiene16; kiesler11; SeeringEtal2019mec). In this new age of generative AI, it is likely that existing moderation strategies will need to adapt in order to remain effective.
@@
This paper qualitatively investigates the experience of moderating AIGC in self-governing communities, in order to explore one dimension of generative AI’s impact on online communities. Not all communities will see AIGC as a concern, but we focus on those that do to illuminate how platform designers and community stewards might support such communities in the face of dramatic technological change (Seering20). How to provide such support is a natural question for the HCI and Social Computing research communities, a central goal of which has been to encourage healthy communities as they strive to meet a variety of individual and collective needs (Cheng17; kiesler11; kraut11; Lampe10).
@@
To this end, we conducted fifteen in-depth, semi-structured interviews with Reddit moderators. We focus on Reddit, a social sharing and news aggregation site, so that we can compare and contrast relatively independent self-governing communities that, nonetheless, share a platform. Reddit is a popular platform to study (proferes21) because of its variety of communities and distributed governance structure: subreddits set and enforce their own rules to supplement site-wide policy (Chandrasekharan18; FieslerEtal2018rrc). Reddit thus allows us to explore how different attitudes and practices towards AIGC organically emerge across communities. We use a qualitative investigation to explore perspectives on what constitutes AIGC and what makes it problematic or challenging. Our goal is to better understand the concerns of moderators in communities that have restricted the use of AIGC, not to provide a representative view of all Reddit moderators or communities. Our work thus explores the following research questions:
@@
RQ1: What are volunteer community moderators’ attitudes towards AIGC in communities that restrict its use?
@@
RQ2: What experiences have these moderators’ had with AIGC in their communities?
@@
Our exploration of RQ1 reveals that our participants restrict the use of AIGC because they are concerned that it may negatively impact the content quality, social dynamics, and governance processes in their communities. Many participants say that AIGC is antithetical to their communities’ purposes and values. Though our participants acknowledge some potential benefits of AIGC, they generally oppose its use and see several specific threats to their communities. Participants say that AIGC is poorly written, inaccurate, and off-topic, and thus threatens to reduce the quality of content in their communities. Our participants also report that AIGC threatens to impact community social dynamics by reducing opportunities for human connection, straining community relationships, and violating shared community values. Finally, AIGC may increase the scale and sophistication of existing types of problematic behavior, making the jobs of moderators more difficult.
@@
Our exploration of RQ2 reveals that even though some participants feel that clarifying rules and norms has rendered AIGC no more threatening than other problematic content, AIGC has increased the workload on already overburdened moderators, as rules about AIGC are hard to enforce. Moderators believe they can detect AIGC, but the process is time-consuming, as they rely on imperfect heuristics that look for certain “tells” in content and behavior. Moderators sometimes use automated detection tools to help with enforcement, but none are reliable enough to fully automate the process. Additionally, moderators acknowledge the sensitivity of accusing a community member of posting AIGC, and emphasize the care that goes into such decisions.
@@
We discuss the implications of these findings for members of online communities, community moderators, and platform designers. We explore the relationship between community values and attitudes towards AIGC and suggest that the arrival of AIGC will drive communities to seek out new ways of demonstrating authenticity. We discuss platform designs that can encourage authenticity and keep members aware of evolving community norms towards AI. We also discuss the potential for bias in moderators’ current detection approaches and explore design solutions that could aid this task by synthesizing various relevant sources of information. Finally, as platforms begin to integrate generative AI features, we suggest that they make these features optional so that communities can retain agency over the types of interactions that they want to encourage.
@@ 2. Related Work
Understanding the challenges posed by AIGC requires an understanding of online communities, their health, and the challenges of moderation. The literature, below, shows that that online communities vary in their reasons for forming and their values; this suggests that community attitudes towards AIGC will be context-dependent. The literature on community health highlights the challenges of dealing with deception, low engagement, and harassment; these might be amplified by AIGC. Finally, research emphasizes the challenging nature of content moderation work; the difficulty of identifying and addressing AIGC might further strain moderators and their relationships with their communities.
@@ 2.1. Online Community Types and Values
Online communities form for different reasons, have different values, and thus may feel threatened by AIGC in different ways. Early work by Preece (preece_empathic_1999) coined the term empathic communities to refer to sites of both emotional and factual communication, where member participation is motivated by a desire for something that professionals cannot provide: authentic, firsthand accounts from others with similar lived experiences. Recent work has shown that these communities can perceive certain types of interactions as inauthentic that are completely acceptable in other contexts (Smith23). We suspect that AI-written responses may be seen as problematic in empathic communities due to their lack of lived experience. Lave and Wegner provide another useful concept, the community of practice (lave1991situated), which denotes a group of people united by a desire to improve their skills by doing, sharing, and receiving feedback from one-another. Since these communities deeply care about process, they might regard contributions as inauthentic if they are made using generative AI rather than their traditional practice. A final useful concept is Jenkins’ idea of knowledge communities in which “members work together to forge new knowledge, often in realms where no traditional expertise exists,”(jenkins_introduction_2006). These communities value accurate knowledge and expertise, which may be threatened by generative AI’s current tendency to produce inaccurate “hallucinations” (augenstein2023factuality). Kraut and Resnick supplement this conceptual literature with practical considerations for platform designers (kraut11) hoping to encourage such communities through systems of trust and reputation (resnick2002trust). These systems of trust may need to evolve to account for a world where contributions made with AI are indistinguishable from those produced by expert community members.
@@
Reddit has been a frequent topic of study for empirical investigations into the different types and values of online communities (proferes21). Reddit’s hierarchical governance structure, in which a set of site-wide policies are supplemented by community-derived and enforced rules, allows it to host many independent communities with distinct purposes and values. This model of community self-moderation (Seering20) allows researchers to study evolving attitudes and practices as they emerge. For example, recent work on the use of governance bots has shown how new governance practices can impact Reddit community dynamics by affecting members’ sense of virtual community (Smith22). Other research has studied community values across the platform through surveys of community members (weld_what_2022), analysis of public subreddit rules (FieslerEtal2018rrc), and logs of of moderation actions (Chandrasekharan18). Findings from these studies emphasize the diverse set of values held by different subreddits. While they identify some commonly occurring rules and values, such as rules against harassment and the importance of trust, they emphasize that rules are largely context-dependent. For example, Fiesler et al (FieslerEtal2018rrc) found that subreddit topic is more predictive than size when predicting the types of rules that a subreddit will have. Recent work introduced the framework of Community Archetypes to describe commonly occurring types of communities on Reddit (Prinster24), which we use to make sense of the variation in community stances towards AIGC. On the whole, the findings from this body of work suggest that Reddit will not have a unified stance towards AIGC and motivates our inquiry into the specific perspectives of moderators from communities with rules governing its use.
@@ 2.2. Generative AI and Threats to Healthy Online Communities
The HCI literature has identified varied threats to online communities which might be amplified by the use of generative AI, such as deception, low engagement, and harassment. Past studies have shown online communities to be vulnerable to strategic influence operations, and recent speculative work has hypothesized that generative AI will increase the volume and believability of such efforts (augenstein2023factuality; goldstein23). Indeed, early empirical studies have found AIGC to be as or more deceptive than human-written content (park_ai_2024) and detected AIGC in fake online reviews (Oak24) and scams (diresta2024spammers). Other studies of intentional deception have focused on the capability of algorithmically controlled social media accounts, or social bots, to deceive unsuspecting users. (Ferrara2016; menczer_addressing_2023). While LLM-powered social bots have already been detected on Twitter (yang23), our work looks beyond social bots to include situations where humans use generative AI tools to create content. We hypothesize that humans posting undisclosed AIGC may be regarded as a subtler form of deception. Findings from AI-Mediated Communication (AIMC) studies suggest that AI affects interpersonal communication (hancock_aimc_2020), causing people to regard others more negatively (rae_effects_2024) and with less trust (Jakesch19) when they perceive that the other is using AI. Past HCI research has explored interface designs that encourage closeness and authenticity by enabling effortful communication, or communication that demonstrates effort and care (Zhang22), between users. It remains to be seen if such design techniques can offset the negative effects that AIGC may have on interpersonal communication.
@@
Another threat to online communities is that of declining participation. The HCI and CSCW literature provides techniques for encouraging community contribution (kraut11) and commitment (ren_encouraging_2012), such as establishing a group identity that allows strong ties to form between members. However, specifically studying how AIGC affects participation in online communities is a relatively new area of research (wei2024understandingimpactaigenerated). Several studies of sites in the Stack Exchange network have demonstrated that the launch of ChatGPT caused an overall decline in website visits, question volume, and number of users, while increasing the complexity of questions asked (burtch_consequences_2024; sanatizadeh_exploring_2023; xue_can_2023). The authors attribute these effects to the relative ease and speed with which users can get simple questions answered by LLM-powered chatbots off-platform. Interestingly, the authors of one study (burtch_consequences_2024) contrast this effect with a measured null-effect on similar Reddit communities, which they suggest are able to avoid a decline in activity because of their emphasis on socialization, rather than pure information exchange. Recent work has also studied the effects of top-down bans on AIGC and found that banning the use of generative AI on Stack Exchange (wang_can_2024) led to a decrease in volume and quality of questions and answers. Our study builds on this line of inquiry by asking if the emergence of AIGC and consequent bans threaten participation in self-governing communities with varied values and needs.
@@
Finally, online communities are often sites of harassment and abuse, a frequent topic in the CSCW literature (cheng_anyone_2017; han_hate_2023); the effects of AI-powered variants, such as malicious synthetic, or deepfake, media (farid_creating_2022; rini_deepfakes_2022) are much less studied. Given the potential for AI-powered abuse, our work explores how community moderators are thinking about these threats and what they are doing to counter them.
@@ 2.3. The Challenges of Volunteer Community Moderators
Subreddits that decide to enact rules about AIGC will turn to already-strained volunteer moderators to enforce these rules (li_all_2022; Wohn19). Moderation involves both “visible” and “invisible” work (Gilbert20), leading researchers to alternatively characterize it as “civic labor” (Matias19), “emotional labor” (Dosono19), and “care work” (gilbert23). The HCI community has a long history of addressing the needs of moderators (Jhaver19), which suggests that research can contribute design solutions specific to AIGC. While existing research has focused on handling harmful content such as misinformation (Bozarth2023), the specific challenges of moderating AIGC have not been studied.
One perpetual challenge for moderators is maintaining alignment between themselves and their community members (Koshy23). Given the power dynamic between moderators and regular members, perceptions of unfair or biased moderation decisions can strain community relationships. Perceptions of bias are not unfounded: existing work has shown that individuals with marginalized identities are more likely to be the target of moderation actions (Haimson21; Thach24). We suspect this will be a problem when moderators enforce rules about AIGC, as research on perception of AIGC has found that humans cannot reliably distinguish between AI and human-generated text (clark21) or images (nightingale22), and use flawed heuristics to make such determinations (Jakesch_2023). Though little is known about moderators’ AI detection practices, recent work has shown that general determinations about AI use are biased according to race and gender stereotypes (mink24). Our study builds on this work by exploring the real-world AI detection practices of community moderators.
The CSCW research community has researched and developed computational tools meant to aid moderators in their work (Chandrasekharan19; Jhaver19; Koshy23; Schluger22; Zhang20). However, detecting AIGC with computational techniques has, so far, proven challenging (yang_survey_2023). While commercial classifiers for AI-generated text claim to be effective (Crothers23), recent research on their practical application questioned their theoretical best-case performance (sadasivan2023). Other work has found these classifiers to perform no better than random chance (weberwulff23), or even worse, to be biased against the writing of non-native English speakers (liang23). This research questions the feasibility of a reliable, general-purpose detection tool that will work on any snippet of text. Still, some studies have had success identifying AIGC in specific types of writing, such as news articles (hanley_machinemade_2024), conference peer reviews (liang_monitoring_2024), and crowd-worker output (veselovsky_prevalence_2023), suggesting that such techniques might be adapted to other contexts. Our study will investigate how moderators think about these tools and currently incorporate them into their workflows.

### v7

@@ Abstract.
Generative AI is altering how we work, learn, communicate, and participate in online communities. How might online communities be changed by generative AI? To start addressing this question, we focused on online community moderators’ experiences with AI-generated content (AIGC). We performed fifteen in-depth, semi-structured interviews with moderators of Reddit communities that restrict the use of AIGC. Our study finds that rules about AIGC are motivated by concerns about content quality, social dynamics, and governance challenges. Moderators fear that, without such rules, AIGC threatens to reduce their communities’ utility and social value. We find that, despite the absence of robust tools for detecting AIGC, moderators were able to somewhat limit the disruption it caused by working with their communities to clarify norms. However, moderators found enforcing AIGC restrictions challenging, as they rely on time-intensive and inaccurate detection heuristics. Our results highlight the importance of supporting community autonomy and self-determination in the face of this sudden technological change, and suggest potential design solutions that may help.
@@ 1. Introduction
Online communities are adjusting their policies and practices to address the proliferation of AI-generated content (AIGC) (Thompson, 2023) produced by generative AI (AI, 2022; Ramesh et al., 2021). The presence of AIGC, whether disclosed or not, may alter the social dynamics of online communities. For example, AI-generated text with subtle inaccuracies may strain collaborative knowledge-sharing communities, such as Stack Exchange (Makyen, 2023) and Wikipedia (Gertner, 2023). Other communities, like Facebook groups where people gather for human connection (Fiesler, 2024), might see AIGC as inauthentic and disruptive. This new form of disruption adds to online communities’ existing challenges, such as harassment or coordinated disinformation campaigns, which generative AI threatens to make cheaper, more effective, and more personalized (Goldstein et al., 2023; Wood, 2024).
@@
As many communities rely on content moderators for stewardship (Grimmelmann, 2015), it is important to understand and document moderators’ firsthand experiences responding to AIGC. Effective moderation can help communities clarify and enforce community norms (Kiene et al., 2016; Kiesler et al., 2011; Seering et al., 2019) in order to stay civil in times of disruptive change. Existing moderation strategies will likely need to adapt in order to remain effective in the face of AIGC. Indeed, there is evidence from recent quantitative research that online communities are enacting rules about the use of AIGC (Lloyd et al., 2025), but exactly how and why they implement such rules requires further study.
@@
In order to more deeply understand generative AI’s potential impact on online communities, this paper qualitatively investigates the experience of moderating AIGC on Reddit. Not all online communities see AIGC as a concern, but we center those that do to illuminate how platform designers and community stewards might help such communities navigate moments of dramatic technological change (Seering, 2020). How to provide such support is a natural question for the HCI and Social Computing research communities, a central goal of which has been to encourage healthy communities as they strive to meet a variety of individual and collective needs (Kiesler et al., 2011; Kraut and Resnick, 2011; Lampe et al., 2010).
@@
To this end, we conducted fifteen semi-structured interviews with volunteer Reddit moderators. We focus on Reddit, a social sharing and news aggregation site, to compare and contrast several relatively independent communities that, nonetheless, share a platform. Reddit is a popular platform to study (Proferes et al., 2021) because of its variety of communities and distributed governance structure: subreddits set and enforce their own rules to supplement site-wide policy (Fiesler et al., 2018). Studying Reddit thus allows us to explore how different attitudes and practices towards AIGC organically emerge across communities. We qualitatively explore moderators’ perspectives on what constitutes AIGC and what makes it problematic. Rather than try to provide a representative view of all Reddit moderators or communities, our goal is to better understand the concerns of moderators in communities that have restricted the use of AIGC. We conducted our interviews within a year of the launch of ChatGPT. While online communities’ concerns about AIGC have continued to evolve (Lloyd et al., 2025), our findings highlight the early concerns of moderators forced to adapt to a new technology. Our work thus explores the following research questions:
@@
RQ1: What are volunteer moderators’ attitudes towards AIGC in online communities that restrict its use?
@@
RQ2: What experiences have these moderators had with AIGC in their communities?
@@
Our exploration of RQ1 reveals that our participants restrict the use of AIGC because they are concerned that it may negatively impact the content quality, social dynamics, and governance processes in their communities. Though our participants acknowledge potential benefits of AIGC, they generally oppose its use—some even see AIGC as completely antithetical to their communities’ purposes and values. Our participants shared several specific threats from AIGC. Participants say that AIGC is poorly written, inaccurate, and off-topic, and thus threatens to reduce the quality of their communities’ content. Participants also report that AIGC may harm community social dynamics by reducing opportunities for human connection, straining relationships, and violating shared values. Finally, participants are concerned that AIGC may increase the scale and sophistication of existing types of problematic behavior, making the jobs of moderators more difficult.
@@
Our exploration of RQ2 reveals that while some participants feel that by clarifying rules and norms they have made AIGC no more threatening than other problematic content, AIGC has increased the workload on already overburdened moderators. Rules about AIGC are hard to enforce: moderators believe they can detect AIGC, but the process relies on time-consuming and imperfect heuristics that look for certain “tells” in content and behavior. Moderators sometimes use automated detection tools to help with enforcement, but none are reliable enough to fully automate the process. Additionally, moderators acknowledge the sensitivity of accusing community members of posting AIGC and emphasize the care that goes into such decisions.
@@
We discuss the implications of these findings for members of online communities, community moderators, and platform designers. We explore the relationship between community values and attitudes towards AIGC and suggest that the arrival of AIGC will drive communities to develop new ways to demonstrate authenticity. We discuss platform designs that can support communities by encouraging authentic communication, as well as others that can keep members aware of evolving community norms. We also discuss potential bias in moderators’ detection methods and explore design solutions that may improve them by synthesizing relevant information sources. Finally, to preserve the types of interactions that communities value, we suggest that platforms allow communities to opt out of any generative AI features.
@@ 2. Related Work
Understanding the challenges posed by AIGC requires an understanding of online communities, their health, and the challenges of moderation. Below we summarize literature showing that online communities vary in their values and their reasons for forming, suggesting that community attitudes towards AIGC will be context-dependent. The literature on community health highlights the challenges of dealing with deception, low engagement, and harassment, all of which might be amplified by AIGC. Finally, we summarize research that emphasizes the challenges of content moderation; the difficulty of identifying and addressing AIGC might further strain moderators and their relationships with their communities.
@@ 2.1. Online Community Types and Values
Online communities form for different reasons, have different values, and thus may feel threatened by AIGC in different ways. Early work by Preece (Preece, 1999) coined the term empathic communities to refer to sites of both emotional and factual communication, where member participation is motivated by a desire for something that professionals cannot provide: authentic, firsthand accounts from others with similar lived experiences. Recent work has shown that these communities consider certain types of interactions inauthentic that are completely acceptable in other contexts (Smith et al., 2023). We suspect that empathic communities may see AI-written posts as problematic due to their lack of lived experience. Another useful concept comes from the work of Lave and Wegner, who use the term community of practice (Lave and Wenger, 1991) to denote a group of people united by a desire to improve their skills by doing, sharing, and receiving feedback from one another. These communities care deeply about process and might regard contributions made with generative AI as inauthentic. Finally, Jenkins’ concept of knowledge communities, in which “members work together to forge new knowledge, often in realms where no traditional expertise exists”(Jenkins, 2006), is very relevant to our study. Knowledge communities value accurate information and expertise, which may be threatened by generative AI’s current tendency to produce inaccurate “hallucinations” (Augenstein et al., 2024). Kraut and Resnick’s work supplements this conceptual literature with practical considerations for platform designers (Kraut and Resnick, 2011), including ways to build systems that support reputation and trust (Resnick and Zeckhauser, 2002). These systems may need to evolve in a world where contributions made with AI are indistinguishable from those produced by experts.
@@
Reddit has been a frequent topic of study for empirical investigations into the different types and values of online communities (Proferes et al., 2021). Reddit’s hierarchical governance structure, in which site-wide policies are supplemented by community-derived and enforced rules, allows it to host many independent communities with distinct purposes and values. This model of community self-moderation (Seering, 2020) allows researchers to study emergent attitudes and practices as they evolve. For example, recent work has shown how the use of governance bots shapes Redditors’ experience of a community, primarily by affecting their sense of virtual community (Smith et al., 2022). Other research has used surveys of community members (Weld et al., 2022), analysis of public subreddit rules (Fiesler et al., 2018; Lloyd et al., 2025), and logs of moderation actions (Chandrasekharan et al., 2018) to study community values across the platform. These studies emphasize the diverse set of values and rules held by different subreddits. Though some are common, such as rules to discourage harassment, many are context-dependent. For example, Fiesler et al (Fiesler et al., 2018) found that subreddit topic is more predictive of subreddits’ rules than size. Recent work introduced the framework of Community Archetypes to describe common types of communities on Reddit (Prinster et al., 2024), which we use to make sense of the variation in community stances toward AIGC. Using this framework, our recent quantitative analysis (Lloyd et al., 2025) shows that Content Generation communities are the most likely to have rules about AI, while Social Support communities are negatively associated with such rules. On the whole, the findings from this literature suggest that subreddits will not have a unified stance towards AIGC; consequently, we seek to understand different moderators’ perspectives.
@@ 2.2. Generative AI and Threats to Healthy Online Communities
The HCI literature has identified varied threats to online communities that generative AI might amplify, such as deception, low engagement, and harassment. Past studies have shown online communities to be vulnerable to strategic influence operations, and recent speculative work has hypothesized that generative AI will increase the volume and believability of such efforts (Augenstein et al., 2024; Goldstein et al., 2023). Indeed, early empirical studies have found AIGC to be as or more deceptive than human-written content (Park et al., 2024), and identified AIGC in fake online reviews (Oak and Shafiq, 2024) and scams (DiResta and Goldstein, 2024). Other studies have focused on the deception capabilities of algorithmically controlled social media accounts, or social bots. (Ferrara et al., 2016; Menczer et al., 2023). While LLM-powered social bots have already been detected on Twitter/X (Yang and Menczer, 2024), our work looks beyond social bots to include situations where humans use generative AI tools to create content. We hypothesize that humans posting undisclosed AIGC may be regarded as a more subtle form of deception. Findings from AI-Mediated Communication (AIMC) studies suggest that AI affects interpersonal communication (Hancock et al., 2020), causing people to regard others more negatively (Rae, 2024; Kadoma et al., 2025) and rate them as less trustworthy (Jakesch et al., 2019) when they perceive that the other is using AI.
@@
Another threat to online communities is that of declining participation. The HCI and CSCW literature provides techniques to encourage community contribution (Kraut and Resnick, 2011) and commitment (Ren et al., 2012), such as by establishing a group identity that can encourage strong ties between members. However, specifically studying how AIGC affects participation in online communities is a relatively new area of research (Wei and Tyson, 2024). Several studies of sites in the Stack Exchange network have demonstrated that the launch of ChatGPT caused an overall decline in website visits, unique users, and question volume, while increasing the complexity of questions asked (Burtch et al., 2024; Sanatizadeh et al., 2025; Xue et al., 2023). The authors attribute these effects to the relative ease and speed with which users can get simple questions answered by LLM-powered chatbots off-platform. Interestingly, the authors of one study (Burtch et al., 2024) contrast this effect with a measured null-effect on similar Reddit communities. The authors suggest that these Reddit communities avoid a decline in activity due to their emphasis on socialization, rather than pure information exchange. Recent work has also studied the effects of top-down bans on AIGC and found that banning the use of generative AI on Stack Exchange (Wang and Zheng, 2024) led to a decrease in volume and quality of questions and answers. Our study builds on this line of inquiry by asking if the emergence of AIGC and consequent bans threaten participation in self-governing communities with varied values and needs.
@@
Finally, online communities are often sites of harassment and abuse, a frequent topic in the CSCW literature (Cheng et al., 2017; Han et al., 2023); the effects of AI-powered variants, such as malicious synthetic, or deepfake, media (Farid, 2022; Rini and Cohen, 2022) are newer and much less studied. Given the potential for AI-powered abuse, our work explores how community moderators are thinking about these threats and what they are doing to counter them.
@@ 2.3. The Challenges of Volunteer Community Moderators
Subreddits that decide to enact rules about AIGC will turn to already-strained volunteer moderators to enforce these rules (Li et al., 2022; Wohn, 2019). Moderation involves both “visible” and “invisible” work (Gilbert, 2020), and is alternatively characterized as “civic labor” (Matias, 2019a), “emotional labor” (Dosono and Semaan, 2019), and “care work” (Gilbert, 2023). The HCI community has a long history of addressing the needs of moderators (Jhaver et al., 2019b). Though existing research has focused on handling harmful content such as misinformation (Bozarth et al., 2023), the specific challenges of moderating AIGC have not been studied.
One perpetual challenge for moderators is maintaining alignment with their community members (Koshy et al., 2023). Given the power dynamic between moderators and regular members, perceptions of unfair or biased moderation decisions can strain community relationships. Perceptions of bias are not unfounded: existing work has shown that individuals with marginalized identities are more likely to be the target of moderation actions (Haimson et al., 2021; Thach et al., 2024). We suspect this will be a problem when moderators enforce rules about AIGC, as research has found that humans cannot reliably distinguish between AI and human-generated text (Clark et al., 2021) or images (Nightingale and Farid, 2022), and use flawed heuristics to make such determinations (Jakesch et al., 2023). Though little is known about moderators’ AI detection practices, recent work has shown that in general, determinations about AI use are biased according to race and gender stereotypes (Mink et al., 2024). Our study builds on this work by exploring the real-world AI detection practices of community moderators.
The CSCW research community has researched and developed computational tools meant to aid moderators in their work (Chandrasekharan et al., 2019; Schluger et al., 2022; Zhang et al., 2020). However, detecting AIGC with computational techniques has, so far, proven challenging (Yang et al., 2024). While commercial classifiers for AI-generated text claim to be effective (Crothers et al., 2023), their performance has been questioned by the research community (Sadasivan et al., 2025). Recent work has found these classifiers to perform no better than random chance (Weber-Wulff et al., 2023), or even worse, to be biased against the writing of non-native English speakers (Liang et al., 2023). This calls into question the feasibility of a reliable, general-purpose detection tool that will work on any snippet of text. Still, other studies have had success identifying AIGC in specific types of writing, such as news articles (Hanley and Durumeric, 2024), conference peer reviews (Liang et al., 2024), crowd-worker output (Veselovsky et al., 2025), and scientific paper abstracts (Kobak et al., 2025), suggesting that such techniques might be adapted to other contexts. Our study will investigate how moderators think about these tools and currently use them in their work.